I appreciate your help on this equation.

2x-5 over 4 = x+7 over 3


thanks for the help

Answer by stanbon(75887)

2x-5 over 4 = x+7 over 3
------------------------------
(2x-5)/4 = (x+7)/3
Cross-multiply to get:
3(2x-5) = 4(x+7)
6x - 15 = 4x + 28
2x = 43
x = 21 1/2
=================
Cheers,
Stan H.